Problem 2-7 Margin Calls (LO3, CFA5) You decide to buy 1700 shares of stock at a price of $94 and an initial margin of 50 percent. What is the maximum percentage decline in the stock price before you will receive a margin call if the maintenance margin is 25 percent? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.) Stock price decline
Answer #1

ANSWER =

Stock Price Decline 33.33%

Explanation :

Amount borrowed = (1,700 * $94)* (1 – .50) = $79,900

Margin call price = ($79,900 / 1,700) / (1 – .25) = $62.67

Stock price decline = ($62.67 – 94) / $94 = -33.33%